Show that if Q ∼ beta(1/2α, 1/2β) then βQ/α(1−Q) ∼ Fα,β.
Answer #1

We are given that Q Beta 22)

Thus, the pdf of Q is given by:

15050: (0 - 1),-30 180 = ()

where B(.,.) denotes the beta funtion.

Now, let

BQ F= 2(1-0) ► aF - aFQ = BQ = (B+aFQ = aF Q = 3 + aF

Thus, we get:

(a)(8 + oF) - (QF)(a) (B+ oF)2 a B (+ F)2

Now, the pdf of F is given by:

SE(P) =se ( p q ft) 2-1 / of + aF af 3-1 aB B+aF) (8 + oF)2 (3 + oF)2 8 + oF) B+F) a 9-1+1 31-1+1 F2-1 B ) (3 + oF) 9–1+3–1+2

9 -1 f(F) - až 33 B(4,1) (B+ aF)***

which is the pdf of \small F_{\alpha,\beta} distribution.

Thus, F-a1-

Hence Proved

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.